A 20 kg crate hangs at the end of a long rope. Find its acceleration when the tension in the rope is (a) 250 N (b) 150 N (c) zero (d) 196 N

1
Expert's answer
2022-03-09T10:32:39-0500

The Newton's second law says

"a=\frac{F}{m}"

(a)

"a=\frac{250}{20}=12.5\:\rm m/s^2"

(b)

"a=\frac{150}{20}=7.5\:\rm m/s^2"

(c)

"a=\frac{0}{20}=0\:\rm m/s^2"

(d)

"a=\frac{196}{20}=9.8\:\rm m/s^2"